if you subtract 13 times a number from three times its square, the result is 10.
Answer by checkley77(12844)

3X^2-13X=10
3X^2-13X-10=0
(3X+2)(X-5)=0
3X+2=0
3X=-2
X=-2/3 ANS.
X-5=0
X=5 ANS.